OpenStudy (anonymous):

Choose the equation of the line passing through the point (–2, 6) and parallel to y = 1/2x – 8. (5 points) y = 1/2x + 7 y = 1/2 x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (mathstudent55):

Given line, y = (1/2)x - 8, has slope 1/2 Parallel lines have the same slope. You want a line that has slope 1/2 and goes through (-2, 6) Use the point-slope form of the equation of a line, where m = slope, and (x1, y1) is point on line. y - y1 = m(x - x1) y - 6 = (1/2)( x - (-2) ) y - 6 = (1/2)x + (1/2)(2) y - 6 = (1/2)x + 1 y = (1/2)x + 7